OpenStudy (anonymous):

Anyone good at algebra two? I am doing conditional probability and cannot figure out how to get the correct answer on a certain question. It says "a box contains three blue marbles, five red marbles, and four white marbles. If one marble is drawn at random, find P(not red l not white). The correct answer is 3/8 but I'm so lost.

OpenStudy (bobo-i-bo):

Given events A and B, \[P(A|B)=\frac{ P(A and B) }{ P(B) }\]

OpenStudy (bobo-i-bo):

In this case, event A is the event of getting not red, event B in the event of getting not white

OpenStudy (bobo-i-bo):

Helpful?

OpenStudy (anonymous):

I would use 7/12 over 8/12 ? is that correct ?

OpenStudy (anonymous):

@bobo-i-bo

OpenStudy (bobo-i-bo):

So you're saying P(A and B)=7/12 and P(B)=8/12?

OpenStudy (anonymous):

yea, I guess so.

OpenStudy (anonymous):

I understand most of this unit, just this one question I don't understand

OpenStudy (bobo-i-bo):

hmmm, so you are wrong for P(A and B). P(A and B)=P(not red and not white) which is the probability of getting any colour but red or white.

OpenStudy (anonymous):

oh oh oh, ok, I think maybe I understand. Let me try it and see

OpenStudy (anonymous):

Omg, i got it, thank you so much! haha

OpenStudy (bobo-i-bo):

:)